wire with resistance3 ohm is  streched from 15. cm to 45 cm..What is the new cross section , new resistance and resistivity

Grade:10

1 Answers

Arun
6 years ago
The volume is constant which means L1*A1 = L2*A2 -----> A2 = (L1/L2)*A1 = A1/3 
R2 = rho*L2/A2 = rho*3L1/A1/3 = 9*rho*L1/A1 
The cross sectional area is 1/3 the inital cross sectional area and the resistance is 9 times the initial resistance = 27Ω